# Constructing g-computation estimators: two case studies in selection bias

### v1

@@ 5 Case Study 2: Confounding and selection bias
Now, consider the joint occurrence of confounding and selection bias described in Zivich et al. and shown in Figure 2 [1]. Note that we recycle our notation (i.e., notation does not carry over from case 1). Let $A$ denote selective serotonin reuptake inhibitor use and $Y$ indicate incident lung cancer in the following 10 years. Again, we are interested in estimated the average causal effect, $\psi$. Let $Z$ denote smoking status, $X$ denote cardiovascular risk score, $U_{1}$ denote depression, $U_{2}$ denote occupational exposures, and $S$ denote if a person completed follow-up. Both $U_{1}$ and $U_{2}$ were unmeasured. As indicated by Figure 2, $Z$ is a confounding variable, and selection bias was related to $X$. Here, there is no superset of variables that addresses both biases. Adjusting for $\{X,Z\}$ opens a backdoor path from $A$ to $Y^{a}$, and only adjusting for $\{X\}$ or $\{Z\}$ does not address both confounding and selection bias. Instead, the parameter of interest can be written as
@@ 5.1 Building an estimator
The identification result in 3 indicates how we an build an iterated g-computation estimator. Starting with the innermost expectation, a model for $E(Y\mid A,Z,X,S=1)$ is estimated. Using that model, we generate pseudo-outcomes under $a$ for all observations, $\hat{Y}^{a}$. Moving to the next expectation, we now fit a model for the pseudo-outcomes given the observed $A$ and $Z$, $E[\hat{Y}^{a}\mid A,Z]$. Regressing pseudo-outcomes is an approach also used by the iterated conditional expectation g-computation algorithm for time-varying confounding [32], a modification of g-computation to address measurement error with partial colliders [33], estimating the conditional average causal effect [34, 35], and estimating the parameters of a marginal structural model [8, 11]. This second model is then used to simulate another set of pseudo-outcomes setting $A$ to $a$, which we denote as $\tilde{Y}^{a}$. The mean of $\tilde{Y}^{a}$ is then our estimate of $\mu^{a}$.
@@
Now, we show that our proposed estimator is consistent and asymptotically normal. Note that the outcome model is unbiased following from maximum likelihood theory. Similarly, unbiasedness of $E\left\{\left[\hat{Y}_{i}^{a}-m(\mathbb{V}_{i};\gamma_{a})\right]\mathbb{V}_{i}^{T}\right\}$ follows from quasi-maximum likelihood [36, 37]. The unbiasedness of the means follows from the identification assumptions and correct model specification. Finally, the estimating equation for the average causal effect follows from the same arguments as before. A more formal proof is provided in Appendix 3.
@@ 5.2 Simulation
Again, we illustrate the proposed estimator with a simulation study. A modified version of the data generating mechanism from Ross et al. was created (see Appendix 3 for details) [33]. Here, we compared a naïve complete-case analysis, standard g-computation adjusting for only $Z$, standard g-computation adjusting for only $X$, standard g-computation adjusting for $\{X,Z\}$, and our proposed iterated g-computation algorithm. The same set of evaluation metrics from case 1 were considered for a simulation of $n:=1000$ with 5000 repetitions. Simulations were conducted in Python 3.9.4 with the same set of packages. Again, we also replicate the results for a single simulated data set in R.
The iterated g-computation estimator had negligible bias and appropriate CI coverage (Table 2), agreeing with the theoretical expectations. All other approaches had bias and below nominal CI coverage.
@@
Bias ESE SER Coverage
Complete-case analysis (naïve) 0.037 0.034 1.01 0%
Standard g-computation, $\{Z\}$ 0.025 0.033 1.01 90%
Standard g-computation, $\{X\}$ 0.042 0.039 1.00 83%
Standard g-computation, $\{X,Z\}$ 0.029 0.038 1.00 90%
Iterated g-computation 0.004 0.036 1.00 95%
@@
ESE: empirical standard error, SER: standard error ratio.
@@ 6 Discussion
Here, we provided two case studies in the context of selection bias on how to develop novel g-computation estimators. These examples illustrate how epidemiologists can translate causal diagrams or identification results into estimators and review some of their properties both theoretically and empirically. The basic principles used here can be extended to build g-computation estimators for more complex causal structures or with additional sources of bias. Furthermore, we found that these basic principles are particularly useful in the scenarios with selection bias, since (1) nonrandom selection (sampling) is often accompanied by missing data on some important variables and then leads to different types of selection bias [38], and (2) researchers tend to rely on naïve complete-case analyses or standard approaches to adjusting for covariates that are related to selection bias, which can result in substantial bias for some nonrandom selection mechanisms.
@@
To develop our estimators and prove some of their theoretical properties, we relied on M-estimation theory. While M-estimators are not required, we find the translation from the discrete steps into estimating functions to be worth the effort given the benefits (e.g., simplified asymptotic proofs, variance estimation via the empirical sandwich variance estimator). However, there are two important limitations to this framework one should be aware of. First, standard M-estimation theory assumes that $\theta$ is of finite-dimension [15]. As such, nonparametric estimators (e.g., Kaplan-Meier) are not immediately ammenable to this framework and other theoretical justifications [39, 40], or structural assumptions are needed [41]. Further, M-estimation assumes that there is a single solution to the estimating equations, which precludes estimators with objective functions with multiple maxima (e.g., neural networks). A third issue is that there are certain restrictions on the structure of the estimating functions, which are reviewed in-depth elsewhere [15]. A notable example of this issue is the Cox proportional hazards model, which can be expressed with estimating equations but is invalid with the standard empirical sandwich variance estimator [42]. Despite these important limitations, M-estimation still covers a broad range of applications in epidemiology.
@@
There are also some challenges to developing estimators that epidemiologists should be aware of. First, identification does not necessarily imply that a parameter is estimable with data, even as the number of observations grows to infinity [43, 44]. This problem arose in the derivation of our estimators, where we also assumed that our parametric model for the outcome process was correctly specified. Second, the proofs of unbiased estimating equations were straightforward for both our estimators. This was mainly due to each estimating function being based on regression models or the mean. However, these proofs may not always be trivial. Therefore, close collaborations with statisticians, biostatisticians, and others experienced in statistical theory are still recommended. Finally, our simulation studies were quite limited. These simulations only verified some general performance metrics in simple scenarios. When developing estimators for a particular use-case, it may be more informative to also develop data generating mechanisms that more closely mimic the use-case [45]. Those simulations may be a better approximation of the expected performance.
@@ Acknowledgments
Research reported in this publication was supported by National Institute of Allergy and Infectious Diseases and the National Institute on Drug Abuse of the National Institutes of Health under award numbers K01AI177102 (PNZ), and K99DA057487 (HL). The content is solely the responsibility of the authors and does not necessarily represent the official views of the National Institutes of Health. Code to replicate the analysis and simulations is available at https://github.com/pzivich/publications-code
@@
$Y$$\sim\begin{cases}\text{Bernoulli}(\text{expit}(0.5+0.75U-1A))&\text{ if }S=1\\ -9999&\text{ if }S=0\\ \end{cases}$
The true value of $\psi$ (as approximated by simulation 10 million observations with $S:=1$ for all observations) was $-0.219$. In the observed data, approximately 20% of participants had missing values of $Y$.
@@
$U_{1}$ $\sim\text{Beroulli}(0.5)$
$U_{2}$ $\sim\text{Beroulli}(0.5)$
$Z$ $\sim\text{Beroulli}(0.5)$
$A$ $\sim\text{Bernoulli}(\text{expit}(-2.3+\log(2)Z+\log(4)U_{1}))$
@@
$S$ $\sim\text{Bernoulli}(\text{expit}(0.25X))$
$Y$$\sim\begin{cases}\text{Bernoulli}(\text{expit}(-2-2A\log(2)Z+\log(2)AZ+\log(4)U_{2}))&\text{ if }S=1\\ -9999&\text{ if }S=0\\ \end{cases}$
The true value of $\psi$ (as approximated by simulation 10 million observations with $S:=1$ for all observations) was $-0.205$. In the observed data, approximately 50% of participants had missing values of $Y$.

### v2

@@ 5 Case Study 2: Confounding and selection bias
Now, consider the joint occurrence of confounding and selection bias described in Zivich et al. and shown in Figure 3 [1]. Note that we recycle our notation (i.e., variable definitions do not carry over from Case 1). Let $A$ denote selective serotonin reuptake inhibitor use and $Y$ indicate incident lung cancer in the following 10 years. Again, we are interested in estimated the average causal effect, $\psi$. Let $Z$ denote smoking status, $X$ denote cardiovascular risk score, $U_{1}$ denote depression, $U_{2}$ denote occupational exposures, and $S$ denote if a person completed follow-up. Both $U_{1}$ and $U_{2}$ were unmeasured. As indicated by Figure 3, $Z$ is a confounding variable, and selection bias was related to $X$. Here, there is no superset of variables that addresses both biases. Adjusting for $\{X,Z\}$ opens a backdoor path from $A$ to $Y^{a}$, and only adjusting for $\{X\}$ or $\{Z\}$ does not address both confounding and selection bias. Instead, the parameter of interest can be written as
@@ 5.1 Building an estimator
The identification result in 3 indicates how we an build an iterated g-computation estimator. Starting with the innermost expectation, a model for $E(Y\mid A,Z,X,S=1)$ is estimated. Using that model, we generate pseudo-outcomes under $a$ for all observations, $\hat{Y}^{a}$. Moving to the next expectation, we now fit a model for the pseudo-outcomes given the observed $A$ and $Z$, $E[\hat{Y}^{a}\mid A,Z]$. Note that the identification results suggest that a model for this value could be separately estimated by levels of $A$ (or equivalently by including interaction terms with $A$ for all covariates in $Z$). However, this need not be the case for the estimator if we assume the corresponding parametric model is correctly specified. In practice, fitting stratified models (or including all the interactions) will likely be preferred due to the weaker modeling assumptions being imposed. Regressing pseudo-outcomes is an approach also used by the iterated conditional expectation g-computation algorithm for time-varying confounding [32], a modification of g-computation to address measurement error with partial colliders [33], estimating the conditional average causal effect [34, 35], and estimating the parameters of a marginal structural model [8, 11]. This second model is then used to simulate another set of pseudo-outcomes setting $A$ to $a$, which we denote as $\tilde{Y}^{a}$. The mean of $\tilde{Y}^{a}$ is then our estimate of $\mu^{a}$.
@@
Now, we argue that our proposed estimator is consistent and asymptotically normal with a more formal proof in Appendix 3. Note that the outcome model is unbiased following from maximum likelihood theory. Similarly, unbiasedness of $E\left\{\left[\hat{Y}_{i}^{a}-m(\mathbb{V}_{i};\gamma_{a})\right]\mathbb{V}_{i}^{T}\right\}$ follows from quasi-maximum likelihood [36, 37]. The unbiasedness of the means follows from the identification assumptions and correct model specification. Finally, the estimating equation for the average causal effect follows from the same arguments as before.
@@ 5.2 Simulation
Again, we verify our proposed estimator using a simulation study. A modified version of the data generating mechanism from Ross et al. was created (see Appendix 3 for details) [33]. Here, we compared a naïve complete-case analysis, standard g-computation adjusting for only $Z$, standard g-computation adjusting for only $X$, standard g-computation adjusting for $\{X,Z\}$, our proposed iterated g-computation algorithm, and an IPW estimator for reference (Appendix 3) [1]. The same set of evaluation metrics from case 1 were considered for a simulation of $n:=1000$ with 5000 repetitions. Simulations were conducted in Python 3.9.4 with the same set of packages. Again, we also replicate the results for a single simulated data set in R.
The iterated g-computation and IPW estimators had negligible bias and appropriate CI coverage (Table 2), agreeing with the theoretical expectations. All other approaches had bias and below nominal CI coverage. The iterated g-computation estimator was the best performing estimator in terms of the RMSE.
@@
Bias ESE RMSE SER Coverage
Complete-case analysis (naïve) 0.037 0.034 0.050 1.01 0%
Standard g-computation, $\{Z\}$ 0.025 0.033 0.041 1.01 90%
Standard g-computation, $\{X\}$ 0.042 0.039 0.058 1.00 83%
Standard g-computation, $\{X,Z\}$ 0.029 0.038 0.047 1.00 90%
Iterated g-computation 0.004 0.036 0.036 1.00 95%
Inverse probability weighting 0.000 0.038 0.038 1.00 95%
@@
ESE: empirical standard error, RMSE: root mean squared error, SER: standard error ratio, CI: confidence interval.
@@ 6 Discussion
Here, we provided two case studies in the context of selection bias on how to develop novel g-computation estimators. These examples illustrate how epidemiologists can translate causal diagrams or identification results into estimators and review some of their properties both theoretically and empirically. The basic principles outlined here can be extended to build g-computation estimators for more complex causal structures or with additional sources of bias. Furthermore, we found that these basic principles are particularly useful in the scenarios with selection bias, since (1) nonrandom selection (sampling) is often accompanied by missing data on some important variables and then leads to different types of selection bias [38], and (2) researchers tend to rely on naïve complete-case analyses or standard approaches to adjusting for covariates that are related to selection bias, which can result in substantial bias for some nonrandom selection mechanisms.
@@
To develop our estimators and prove some of their theoretical properties, we relied on M-estimation theory. While M-estimators are not required, we find the additional step of translating from discrete steps into estimating functions to be worth the additional effort (e.g., simplified asymptotic proofs, variance estimation). However, there are important limitations to this framework one should be aware of. First, standard M-estimation theory assumes that $\theta$ is of finite-dimension [15]. As such, nonparametric estimators (e.g., Kaplan-Meier) are not immediately ammenable to this framework. Other theoretical justifications [39, 40], or structural assumptions are needed [41]. Further, M-estimation assumes that there is a single solution to the estimating equations, which precludes estimators with objective functions with multiple maxima (e.g., neural networks). A third issue is that there are certain restrictions on the structure of the estimating functions, which are reviewed in-depth elsewhere [15]. Notables examples that prevent use of the standard empirical sandwich variance estimator are the Cox proportional hazards model and the L1 penalty (i.e., LASSO) [42, 43, 44]. Fourth, all relevant estimation steps must be included in the stacked estimating functions for valid inference with the empirical sandwich variance estimator. Importantly, this included model or variable selection steps that are not based on subject matter knowledge. While the L1 penalty presents issues for inference, the L2 penalty or smoothed versions of the L1 penalty provide ways forward [43, 44, 45]. Despite these important limitations, M-estimation still covers a broad range of applications in epidemiology.
@@
There are also some challenges to developing estimators that epidemiologists should be aware of. First, identification does not necessarily imply that a parameter is estimable with data, even as the number of observations grows to infinity [46, 47]. This problem arose in the derivation of our estimators, where we also assumed that our parametric model for the outcome process was correctly specified. Second, the proofs of unbiased estimating equations were straightforward for both our estimators. This was mainly due to each estimating function being based on regression models or the mean. However, these proofs may not always be trivial. Therefore, close collaborations with statisticians, biostatisticians, and others experienced in statistical theory are still recommended.
There are several directions in which this work could be further developed. First, one might consider how models for different processes can be combined to construct augmented inverse probability weighting (AIPW) or targeted maximum likelihood estimators [48, 49], which may offer robustness to different modeling assumptions. One approach to deriving an AIPW estimator is to derive the influence function from the identification results, which can be done using several different methods [50, 51]. Given the connection between estimating function and influence functions [52], that AIPW estimator could then be expressed with estimating functions to possibly allow for multiply robust point and variance estimation with parametric models [53]. There also might be interest in using flexible machine learning algorithms instead of parametric regression models. While the use of machine learning for estimation presents certain challenges, these can often be addressed through the use of AIPW estimators [10, 54, 55, 56]. Again, it is helpful to turn to influence functions here. Influence functions provide ways to study additional asymptotic properties of estimators, including convergence rates with flexible machine learning algorithms [51]. While the influence function approach can provide additional results, g-computation estimators built with estimating functions can be directly developed from identification results and thus are a useful first step. Next, our examples did not include the setting where outcomes influence selection (e.g., case-control studies). In these settings, identifying average cusal effects in the entire sample might be challenging. Ultimately, the recoverability of causal effects in the presence of selection bias will depend on the underlying causal structure, the information available, and the parameter of interest. While the proposed g-computation estimators may no longer apply, the underlying ideas of building up estimators from identification results will. Finally, the aim of our simulations was to verify the derivations and showcase some general finite sample performance, so selection bias in our mechanisms was quite substantial. Exploring the properties of all the competing estimators under different selection mechanisms with varying relationship strengths and sample sizes or mimicing particular use-cases in future work would be informative to guide their practical use [57].
@@ Acknowledgments
Research reported in this publication was supported by National Institute of Allergy and Infectious Diseases and the National Institute on Drug Abuse of the National Institutes of Health under award numbers K01AI177102 (PNZ), and K99DA057487 (HL). The content is solely the responsibility of the authors and does not necessarily represent the official views of the National Institutes of Health. Python and R code to replicate the analysis and simulations is available at https://github.com/pzivich/publications-code
@@
[…]
The true value of $\psi$ (as approximated by simulation 10 million observations with $S:=1$ for all observations) was 0.219-0.219- 0.219. In the observed data, approximately 20% of participants had missing values of $Y$.
A2.4: Inverse Probability Weighting Estimator
The corresponding IPW expression for the identification results in A2.1 is
$\mu^{a}=E\left[Y\times I(A=a)\times\frac{S}{\Pr(S=1\mid X)}\right]$
and the corresponding estimating functions for the Hajek IPW are
[$\begin{bmatrix}\left\{S_{i}-\text{expit}(\mathbb{S}_{i}\hat{\alpha}^{T})\right\}\mathbb{S}_{i}^{T}\\ \frac{S_{i}}{\text{expit}(\mathbb{S}_{i}\hat{\alpha}^{T})}A_{i}(Y_{i}-\hat{\mu}^{1})\\ \frac{S_{i}}{\text{expit}(\mathbb{S}_{i}\hat{\alpha}^{T})}\left\{1-A_{i}\right\}(Y_{i}-\hat{\mu}^{0})\\ (\hat{\mu}^{1}-\hat{\mu}^{0})-\hat{\psi}\end{bmatrix}$
where $\text{expit}(\mathbb{S}_{i}\hat{\alpha}^{T})$ is a logistic regression model for the probability of $S$ given $X$ (i.e., $\mathbb{S}$ is a design matrix that includes $X$, $\text{expit}(x)=\{1+\exp(-x)\}^{-1}$). Note that we do not provide a proof of the asymptotic properties of this estimators as it is beyond the scope of this paper.
@@
[…]
@@
[…]
The true value of $\psi$ (as approximated by simulation 10 million observations with $S:=1$ for all observations) was 0.205-0.205- 0.205. In the observed data, approximately 50% of participants had missing values of $Y$.
A2.4: Inverse Probability Weighting Estimator
The corresponding IPW expression for the identification results in A3.1 is
$\mu^{a}=E\left[Y\times\frac{I(A=a)}{\Pr(A=a\mid Z)}\times\frac{S}{\Pr(S=1\mid X)}\right]$
and the corresponding estimating functions for the Hajek IPW are
[$\begin{bmatrix}\left\{S_{i}-\text{expit}(\mathbb{S}_{i}\hat{\alpha}^{T})\right\}\mathbb{S}_{i}^{T}\\ \left\{A_{i}-\text{expit}(\mathbb{A}_{i}\hat{\delta}^{T})\right\}\mathbb{A}_{i}^{T}\\ \frac{S_{i}}{\text{expit}(\mathbb{S}_{i}\hat{\alpha}^{T})}\times\frac{A_{i}}{\text{expit}(\mathbb{A}_{i}\hat{\delta}^{T})}\times(Y_{i}-\hat{\mu}^{1})\\ \frac{S_{i}}{\text{expit}(\mathbb{S}_{i}\hat{\alpha}^{T})}\times\frac{1-A_{i}}{1-\text{expit}(\mathbb{A}_{i}\hat{\delta}^{T})}\times(Y_{i}-\hat{\mu}^{0})\\ (\hat{\mu}^{1}-\hat{\mu}^{0})-\hat{\psi}\end{bmatrix}$
where $\text{expit}(\mathbb{S}_{i}\hat{\alpha}^{T})$ is a logistic regression model for the probability of $S$ given $X$ (i.e., $\mathbb{S}$ is a design matrix that includes $X$) and $\text{expit}(\mathbb{A}_{i}\hat{\delta}^{T})$ is a logistic regression model for the probability of $A$ given $Z$ (i.e., $\mathbb{A}$ is a design matrix that includes $Z$). Note that we do not provide a proof of the asymptotic properties of this estimators as it is beyond the scope of this paper.